In a particular state, license plates must consist of 3 letters followed by 3 numbers. How many such plates exist if the first letter cannot be a z and the plate contains at least one 4?
What threw me off is the at least one 4 part.

Answer by stanbon(57347)

In a particular state, license plates must consist of 3 letters followed by 3 numbers. How many such plates exist if the first letter cannot be a z and the plate contains at least one 4?
--------------------------------
# of plates = # possible - # with no 4's
----
# possible = 25*26^2*10^3 = 16900000
# with no 4's = 25*26^2*9^3 = 12320100
----
# of plates = 16900000-12320100 = 4579900
=================
Cheers,
Stan H.
==============